46. Storm, Inc. purchased the following available-for-sale securities during Year 1, its first year of operations: Na...

46.

Storm, Inc. purchased the following available-for-sale securities during Year 1, its first year of operations:

Name Number of Shares Cost
Dust Devil, Inc. 1,880 $80,840
Gale Co. 800 64,000
Whirlwind Co. 2,850 111,150
Total $255,990

The market price per share for the available-for-sale security portfolio on December 31, Year 1, was as follows:

Market Price per Share,

Dec. 31, Year 1

Dust Devil, Inc. $39
Gale Co. 75
Whirlwind Co. 41

Required:

A. Provide the journal entry to adjust the available-for-sale security portfolio to fair value on December 31, Year 1. Refer to the Chart of Accounts for exact wording of account titles.
B. Describe the income statement impact from the December 31, Year 1, journal entry.

A. Provide the journal entry to adjust the available-for-sale security portfolio to fair value on December 31, Year 1. Refer to the Chart of Accounts for exact wording of account titles.

PAGE 10

JOURNAL

ACCOUNTING EQUATION

DATE DESCRIPTION POST. REF. DEBIT CREDIT ASSETS LIABILITIES EQUITY

1

Adjusting Entries

2

3

B. Describe the income statement impact from the December 31, Year 1, journal entry.

There --------- income statement impact from the December 31, Year 1, adjusting entry. Unrealized Gain (Loss) on Available-for-Sale Investments is reported in the -------------- .

On December 31, Year 1, Unrealized Gain or Loss on Available-for-Sale Investments would be disclosed as follows (if a loss, use a minus sign): ---------------

Unrealized gain (loss) on available-for-sale investments:
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Thanks!
Storm, Inc.
Calculation of fair value of investment
Shares Number Price per share Fair value
Dust Devil, Inc.                                     1,880.00                 39.00     73,320.00
Gale Co.                                        800.00                 75.00     60,000.00
Whirlwind Co.                                     2,850.00                 41.00 116,850.00
Answer a
Schedule of cost and fair value of the portfolio at the end of Dec 31, Year 1.
//img.homeworklib.com/questions/b2883e30-12c0-11ea-98b1-fb5b0a3d66ab.gif?x-oss-process=image/resize,w_560
Investment
Cost Fair value Unrealized Gain/(Loss)
Dust Devil, Inc.                                  80,840.00         73,320.00      (7,520.00)
Gale Co.                                  64,000.00         60,000.00      (4,000.00)
Whirlwind Co.                                111,150.00       116,850.00       5,700.00
Net Impact     (5,820.00)
Answer b Journal Entry
//img.homeworklib.com/questions/b2883e30-12c0-11ea-98b1-fb5b0a3d66ab.gif?x-oss-process=image/resize,w_560
Date
Account Debit $ Credit $ It will impact Equity.
Dec 31 Unrealized Loss            7,520.00
Investment in Dust Devil, Inc.       7,520.00
//img.homeworklib.com/questions/b2883e30-12c0-11ea-98b1-fb5b0a3d66ab.gif?x-oss-process=image/resize,w_560
Dec 31
Unrealized Loss            4,000.00
Investment in Gale Co.       4,000.00
//img.homeworklib.com/questions/b2883e30-12c0-11ea-98b1-fb5b0a3d66ab.gif?x-oss-process=image/resize,w_560
Dec 31
Investment in Whirlwind Co.            5,700.00
//img.homeworklib.com/questions/b2883e30-12c0-11ea-98b1-fb5b0a3d66ab.gif?x-oss-process=image/resize,w_560
Unrealized Gain       5,700.00
Unrealized loss on available-for-sale investments: $ 5,820.